OpenStudy (bossimbacon):

What is the x-intercept of the line with the equation ? 1/3x+y=-15 ( , 0) @shahd-owies

OpenStudy (anonymous):

1/3x+y=-15 to find the x intersection that means to find x when y=0 now substitute y=0 in the equation 1/3x+y=-15 1/3x+0=-15 1/3x=-15 x=-15*3 x=-45 so the x intersection is when x=-45 (-45,0)